OpenStudy (anonymous):

How do I find the sum of the angle measures of a polygon? Can I have an example please.

Directrix (directrix):

The sum of the interior angle measures of a convex polygon of n sides is the following: (n - 2) * 180

Directrix (directrix):

Example: Find the sum of the interior angle measures of a convex pentagon. A pentagon has 5 sides. The sum of the interior angles is (5 - 2) * 180 = 3 * 180 = 540.

OpenStudy (anonymous):

The sum of the exterior angle of a polygon is 360